OpenStudy (anonymous):

The endpoints of a line segment are at the coordinates (3, -4, 6) and (5, 7, -2). What is the midpoint of the segment?

OpenStudy (anonymous):

*You asked this question twice so I'm just pasting again.* To solve this, all you have to do is divide the x coordinates together, the y coordinates together, and the z coordinates together then divide by 2. So we have (3, -4, 6) and (5, 7, -2), now let's solve it: (3 + 5) ÷ 2 = 4 (-4 + 7) ÷ 2 = 1.5 (6 + -2) ÷ 2 = 2 So your answer is A. (4, 1.5, 2)